# Human-In-The-Loop Machine Learning for Safe and Ethical Autonomous Vehicles: Principles, Challenges, and Opportunities

### v1

@@ Abstract
Rapid advances in Machine Learning (ML) have triggered new trends in Autonomous Vehicles (AVs). ML algorithms play a crucial role in interpreting sensor data, predicting potential hazards, and optimizing navigation strategies. However, achieving full autonomy in cluttered and complex situations, such as intricate intersections, diverse sceneries, varied trajectories, and complex missions, is still challenging, and the cost of data labeling remains a significant bottleneck. The adaptability and robustness of humans in complex scenarios motivate the inclusion of humans in ML process, leveraging their creativity, ethical power, and emotional intelligence to improve ML effectiveness. The scientific community knows this approach as Human-In-The-Loop Machine Learning (HITL-ML). Towards safe and ethical autonomy, we present a review of HITL-ML for AVs, focusing on Curriculum Learning (CL), Human-In-The-Loop Reinforcement Learning (HITL-RL), Active Learning (AL), and ethical principles. In CL, human experts systematically train ML models by starting with simple tasks and gradually progressing to more difficult ones. HITL-RL significantly enhances the RL process by incorporating human input through techniques like reward shaping, action injection, and interactive learning. AL streamlines the annotation process by targeting specific instances that need to be labeled with human oversight, reducing the overall time and cost associated with training. Ethical principles must be embedded in AVs to align their behavior with societal values and norms. In addition, we provide insights and specify future research directions.
@@
Human-In-The-Loop, Machine Learning, Curriculum Learning, Reinforcement Learning, Active Learning, Ethical, Autonomous Vehicles, Unmanned Aerial Vehicles
@@ I Introduction
Recent developments in Autonomous Vehicles (AVs) have been useful in improving efficiency and safety[1]. AVs have ushered in a substantial transformation in terms of automation and connectivity. The global AVs market size was valued at USD 1,500.3\$ billion in 2022 and is projected to grow from USD 1,921.1\$ billion in 2023 to USD 13,632.4\$ billion by 2030[2]. Commercially envisioned and available self-driving cars can perform functions such as lane-changing, highway driving, and autonomous parking by sensing the environment using built-in technologies[3]. Ground-based autonomous systems can be paired with aerial ones to drive a significant advancement in transportation and technology and offer a range of exciting possibilities, including surveillance, monitoring, traffic management, remote sensing, and autonomous control [4], [5]. Current trends in AVs are leveraging Machine Learning (ML)[6], tightening security [7], [8], and facilitating data and command dissemination [9], [10] to decrease accidents, fatalities, injuries, pollution, and traffic congestion while fostering a more equitable transportation infrastructure.
To make it easier to understand the terminology used in this paper a list of all acronyms can be found in Table I.
Figure 1: The workflow of HITL-ML for AVs includes several key steps: developing an ML model for decision-making and predictive modeling using PPO, applying human validation and annotation through HITL-RL and AL to improve safety and foster public trust, conducting model training where data is organized and refined with CL to enhance learning efficiency and performance, and real-world deployment while considering ethical implications.
@@ I-A Motivation and Background
Human-In-The-Loop Machine Learning (HITL-ML) combines human creativity, ethical judgment, and emotional intelligence with the learning power of ML algorithms. For the foreseeable future, successfully deploying ML in AVs will require close collaboration and integration with humans. Human intuition and judgment help AVs encounter situations that are not well-represented in training data. Decision making for ethical dilemmas in unavoidable accidents, require human-like reasoning that current ML models do not possess. Feedback from human operators can improve ML models and expedite gaining public trust in AVs. Humans can be incorporated into the foundations of ML through collaboration, and oversight. In collaboration, humans and ML models are near equals in performance but not in abilities. Human excels at capturing region-specific behaviors and cultural norms, providing context and insights that ML algorithms might miss. ML model can ask for help when they are either unconfident in their performance and/or think the human can better carry out the task.
Consider an AV navigating through a city and dealing with a complex driving scenario, e.g. a chaotic intersection in a foreign country, where the traffic rules differ significantly from those on which the AV has been trained. The AV’s ML model initially relies on its standard environmental rewards, such as staying in its lane, obeying traffic signs and avoiding obstacles. However, in this unfamiliar situation, the model’s performance begins to degrade, in this case it can request human assistance to refine its understanding of the situation. A human operator, aware of local driving habits and cultural nuances, could then supplement the ML model’s environmental rewards with human-defined rewards. For example,the human might adjust the reward structure to prioritize yielding to pedestrians who are likely to cross the street outside of designated crosswalks, a behavior common in that region but not explicitly covered by the original training data.
@@
AVs still encounter issues that require human oversight. ML models are strictly better than humans at tasks such as predictive analytics and risk assessment, but we still need to monitor them to ensure safety and alignment with our goals (oversight). Tesla’s Autopilot, for instance, has problems when driving on local roads in the presence of people and relies on the driver’s supervision to navigate effectively [11]. Google’s self-driving cars still make mistakes, such as being confused by fixed-gear cyclists at stop signs [12]. In October 2023, Cruise’s permit to test AVs in California was suspended after a serious accident in which a car dragged a woman across the road. Waymo recently recalled some of its vehicles after two collided with the same tow truck within minutes [13]. This ongoing need for human involvement underscores why AVs are still several years away from being ready for widespread consumer use [14]. As demonstrated by the recent challenges faced by companies like Tesla and Waymo, the effective deployment of AVs requires a hybrid approach where human oversight continues to play a crucial role in ensuring safety and reliability.
@@
As depicted in Fig. 1, the workflow of HITL-ML for AVs comprises several key steps, developing an ML model using predictive algorithms, applying human validation and annotation with expert oversight, conducting model training where input data is organized and refined, and real-world deployment while considering ethical implications. Predictive algorithms such as Proximal Policy Optimization (PPO) are initially used, with human experts monitoring their functionalities where they replace dangerous actions in critical situations and modify or supplement the reward function. Moreover, Active Learning (AL) can detect instances where AVs lack adequate confidence and refer them to human annotators[15]. The model is trained after human validation, annotation, and feedback from human experts. Meanwhile, Curriculum Learning (CL) with the knowledge of human experts provides sorted and organized input for AVs[16], facilitating and expediting the learning process. Finally, the model is deployed in the real world, considering ethical implications.
Rapid progress in user interface design[17], regulatory frameworks [18], AV security[19],[20] and explainable AI [21] have significantly improved the practicality of HITL approaches for AVs. Improved user interfaces provide intuitive controls and real-time feedback, while regulatory requirements ensure human oversight for safety and ethical compliance. Developments in AV security safeguard against cyber threats, and research in explainable AI enhances the transparency of decision-making processes, making it easier for human operators to understand and interact with AVs. Together, these advancements enable effective human oversight and intervention, crucial for the safe and ethical operation of AVs.
Figure 2: The diagram of CL for AVs highlights a process where tasks are divided into smaller, manageable parts, progressing from easy to more difficult challenges. Data associated with each part is sequentially fed to the model to enhance learning efficiency and improve overall performance.
@@ I-B Existing Surveys and Tutorials
In the literature, a few works consider HITL-ML. An up-to-date overview of the latest results on understanding and utilizing the human factor in different contexts of mobile networks is provided by the authors of [22]. Expanding on HITL-ML strategies, [23] presents a comprehensive review of state-of-the-art techniques, highlighting different strategies for incorporating humans into the ML process. From a data perspective, [24] surveys HITL-ML, addressing how human interaction can enhance data quality and model performance. The authors of [25] review the main concepts and methods for HITL learning, paving the way for developing frameworks that facilitate human-machine teaming through safe learning and anomaly prediction. Methodologies, challenges, and opportunities associated with real-world applications of HITL systems are reviewed by [26], discussing key issues like data quality, bias, and user engagement. [27] reviews the literature on HITL, emphasizing the role of human interaction in enhancing computer vision models’ performance and accuracy.
Specifically, in the realm of CL within RL, [28] provides a framework and an overview of existing CL methods. The authors classify these methods based on their assumptions, capabilities, and goals, offering insights into how CL can systematically improve RL performance. [29] examines recent advances in AI, ML, and DL in advanced robotic systems, and discusses their applications in robot modification. [30] reviews the usage of AI for AVs development, and elaborates on critical issues such as functional needs and challenges in implementing AI-driven solutions for AVs. From a computer vision system design perspective, [31] presents a review of HITL-Cyber-Physical Systems (CPSs), providing insights into how human factors can be integrated into CPSs to enhance their functionality and reliability. AD technologies have seen substantial advancements, driven by developments in ML and AI. The authors of [32] review the End-to-End AD stack, categorizing automated driving tasks using neural networks in an End-to-End manner. Surveying essential components and cutting-edge technologies, [33] explores the development and implementation of AD systems. This survey highlights the key technologies that enable AD and the challenges associated with their deployment. [34] and [35] address ethical dilemmas associated with different aspects of AVs.
@@ I-C Contribution
In this paper, we review HITL-ML to support the safe and ethical operation of AVs, facilitate the management of complex and unpredictable scenarios from busy local roads to intricate intersections, trajectories, and missions improve their trustworthiness, and enhance public confidence in them. Our paper differs from the literature by emphasizing the critical role of HITL-ML in ensuring the safe and ethical operation of AVs. We focus on CL, HITL-RL, AL , and ethical principles tailored to the unique challenges and opportunities in AVs. The key contributions of this paper are listed as follows:
Figure 3: Schematic diagram of content and organization.
  • We provide a comprehensive overview of the application of CL in AVs and show how effective it is to improve generalization and convergence speed by structuring the training process from simpler to more complex tasks. We highlight the synergy between CL and RL and show how their combination can reduce training time and improve AV performance on tasks such as navigation, path planning and obstacle avoidance.
  • We provide an in-depth overview of the latest HITL-RL methods and their applications, highlighting how human involvement can enhance the RL process through techniques such as reward shaping, action injection and interactive learning. We also present detailed examples from recent studies showing various HITL-RL frameworks that utilize human feedback to refine AV learning processes and ensure real-time safety interventions.
@@ I-D Paper Structure and Organization
For the convenience of readers, Fig. 3 shows the organization of this paper. In Section II, we introduce CL and its significant role in improving learning speed in AVs, review the literature, and conclude with key findings and insights. In Section III, we introduce HITL-RL, review the literature, and conclude with key findings and insights. In Section IV, we discuss AL and examine its role in facilitating data annotation in AVs, review the literature, and conclude with key findings and insights. In Section V, we review the literature on the ethical aspects of AVs, provide guidelines to achieve ethical AVs using HITL-ML, and conclude with key findings and insights. In Section VI, we highlight future research directions. Finally, we conclude the paper in Section VII.
@@
[36] To assist the UAV in obstacle avoidance. DRL Navigation / UAV
[37] To improve the agent’s learning ability and performance. DRL Intelligence / AV
[38] To tackle sparse reward and low learning efficiency. RL Exploration / AUV
[39] To facilitate mapless obstacle avoidance. DRL Navigation / UAV swarm
[40] To improve the learning performance of RL DRL AD
[41] To mitigate in-flight UAV anomalies by using semi-supervised SVM AL Anomaly detection / UAV
[42] To develop a new path planning framework for AL AL Path planning / Semantic mapping
[43] To detect unknown network attacks in the UAV network. AL IDS / UAV
[15] To facilitate manual labeling of large-scale surveillance data. AL VTR / ITS
[44] To learn the decentralized flocking and obstacle avoidance policy CL Obstacle avoidance / UAV swarm
[45] To reduce the required training time and improve agent performance CL AD
[46] To enable mission planning for controlling a UAV ML Mission planning / UAV
[47] To address 3D object detection in AV AL Object detection / UAV
[48] To study collision-free flocking control problem for scalable and large-scale fixed-wing UAV CL Flocking problem / UAV swarm
[49] To learn an end-to-end urban driving policy for CARLA driving simulator CL Driving simulator
@@ II Curriculum Learning
This section reviews recent studies that applied CL to AVs. CL is an ML technique that involves the systematic training of models by first focusing on simple tasks and gradually progressing to more difficult and complex tasks, with human experts guiding the design and adjustment of the learning curriculum to ensure optimal learning outcomes. This technique offers the advantages of generalization and a fast convergence speed [50]. Table II provides a comparative summary of recent research papers exploring HITL- ML comprising CL, HITL-RL, and AL.
II-A Preliminaries
@@
CL is an ML technique that structures the training process by starting with simpler tasks and gradually progressing to more complex ones. This approach enhances generalization and convergence speed, making it particularly effective for training AVs. Recent studies have applied CL to AVs in different ways, including curriculum-based multi-agent DRL for decentralized flocking and obstacle avoidance, collision-free flocking control in large-scale UAV swarms, and adaptive curriculum strategies for efficient target search in disaster scenarios using UAV swarms. These applications demonstrate the effectiveness of CL in guiding autonomous agents through progressively challenging tasks to achieve optimal performance.
@@
As depicted in Fig. 2, the concept of CL for AVs is highlighted. Tasks are divided into smaller, manageable parts, progressing from easy to more difficult challenges. This method allows the AVs to build foundational skills with simpler tasks before gradually tackling more complex scenarios, thereby improving learning efficiency and performance over time. The structured progression ensures that the vehicle can handle increasingly difficult situations with greater accuracy and reliability. In the realm of AVs, the combination of CL and RL can reduce the required training time, improve agent performance, and help learn better policies [45]. AVs require mastery of complex tasks such as navigation, path planning, and obstacle avoidance. CL can aid the learning process by ordering these tasks from simple to complex.
II-B Literature Review
II-B1 Navigation and Data Collection

### v2

@@ Abstract
Rapid advances in Machine Learning (ML) have triggered new trends in Autonomous Vehicles (AVs). ML algorithms play a crucial role in interpreting sensor data, predicting potential hazards, and optimizing navigation strategies. However, achieving full autonomy in cluttered and complex situations, such as intricate intersections, diverse sceneries, varied trajectories, and complex missions, is still challenging, and the cost of data labeling remains a significant bottleneck. The adaptability and robustness of humans in complex scenarios motivate the inclusion of humans in the ML process, leveraging their creativity, ethical power, and emotional intelligence to improve ML effectiveness. The scientific community knows this approach as Human-In-The-Loop Machine Learning (HITL-ML). Towards safe and ethical autonomy, we present a review of HITL-ML for AVs, focusing on Curriculum Learning (CL), Human-In-The-Loop Reinforcement Learning (HITL-RL), Active Learning (AL), and ethical principles. In CL, human experts systematically train ML models by starting with simple tasks and gradually progressing to more difficult ones. HITL-RL significantly enhances the RL process by incorporating human input through techniques like reward shaping, action injection, and interactive learning. AL streamlines the annotation process by targeting specific instances that need to be labeled with human oversight, reducing the overall time and cost associated with training. Ethical principles must be embedded in AVs to align their behavior with societal values and norms. In addition, we provide insights and specify future research directions.
@@
Human-In-The-Loop, Machine Learning, Curriculum Learning, Reinforcement Learning, Active Learning, Ethical Principles, Autonomous Vehicles, Unmanned Aerial Vehicles
@@ I Introduction
Recent developments in Autonomous Vehicles (AVs) have been useful in improving efficiency and safety[1]. AVs have ushered in a substantial transformation in terms of automation and connectivity. The global AVs market size was valued at USD 1,500.3\$ billion in 2022 and is projected to grow from USD 1,921.1\$ billion in 2023 to USD 13,632.4\$ billion by 2030[2]. Commercially envisioned and available self-driving cars can perform functions such as lane-changing, highway driving, and autonomous parking by sensing the environment using built-in technologies[3]. Ground-based autonomous systems can be paired with aerial ones to drive a significant advancement in transportation and technology and offer a range of exciting possibilities, including surveillance, monitoring, traffic management, remote sensing, and autonomous control [4][5]. Current trends in AVs include leveraging Machine Learning (ML)[6], tightening security [7][8], and facilitating data and command dissemination [9][10] to decrease accidents, fatalities, injuries, pollution, and traffic congestion while fostering a more equitable transportation infrastructure.
Figure 1: The workflow of HITL-ML for AVs includes several key steps: 1) developing an ML model for decision-making and predictive modeling using PPO, 2) applying human validation and annotation through HITL-RL and AL to improve safety and foster public trust, 3) conducting model training where data is organized and refined with CL to enhance learning efficiency and performance, and 4) real-world deployment while considering ethical implications.
ML is essential for the operation of AVs because it enables AVs to learn from large datasets and create predictive models for their environment. ML algorithms facilitate the detection and classification of objects, pedestrians and other vehicles, as well as the prediction of their future trajectories. These predictive insights are critical to enabling AVs to make real-time braking, turning, acceleration and steering decisions. The use of AVs fundamentally depends on ML-driven systems to perform key functions such as perception[11], state estimation[12], path planning[13], traffic rule compliance and decision making in dynamic and uncertain environments.
Table I presents a list of the acronyms used throughout this paper for easy reference.
@@ I-A Motivation and Background
Human-In-The-Loop Machine Learning (HITL-ML) combines human creativity, ethical judgment, and emotional intelligence with the learning power of ML algorithms. Human intuition and judgment help AVs handle situations that are not well-represented in training data [14]. Decision making for ethical dilemmas in unavoidable accidents requires human-like reasoning that current ML models do not possess. Human-like reasoning is crucial in these scenarios because it incorporates context, societal norms, and a nuanced understanding of moral principles, capacities that current ML models lack. Human input and guidance can enhance ML models and accelerate the process of building public trust in AVs. Humans can be incorporated into the foundations of ML through collaboration, and oversight. In collaboration, humans and ML models are near equals in performance but not in abilities. Humans excel at capturing region-specific behaviors and cultural norms, providing context and insights that ML algorithms might miss. ML models can ask for help when they are either unconfident in their performance and/or think humans can better carry out the task.
Imagine an AV driving on a local road where pedestrians frequently cross unpredictably and outside of designated crosswalks. In such a scenario, the AV’s ML model may struggle to adapt, resulting in dangerous actions such as abrupt stops or failure to yield appropriately. When the AV encounters these challenges, it can request human intervention to prevent dangerous behaviors and refine its response. A human operator familiar with the local conditions can override or adjust the model’s actions, supplementing the reward system to prioritize pedestrian safety, even in situations that deviate from the AV’s original training data. This human input ensures that the AV can navigate the environment safely and effectively despite the limitations of its original training.
@@
ML models are strictly better than humans at tasks such as predictive analytics and risk assessment, but we still need to monitor them to ensure safety and alignment with our goals (oversight). Tesla’s Autopilot, for instance, has problems when driving on local roads in the presence of people and relies on the driver’s supervision to navigate effectively [15]. Google’s self-driving cars still make mistakes, such as being confused by fixed-gear cyclists at stop signs [16]. In October 2023, Cruise’s permit to test AVs in California was suspended after a serious accident in which a car dragged a woman across the road. Waymo recently recalled some of its vehicles after two collided with the same tow truck within minutes [17]. This ongoing need for human involvement underscores why AVs are still several years away from being ready for widespread consumer use [18]. As demonstrated by the recent challenges faced by companies like Tesla and Waymo, the effective deployment of AVs requires a hybrid approach where human oversight continues to play a crucial role in ensuring safety and reliability.
@@
As shown in Fig. 1, the workflow of HITL-ML for AVs comprises several key steps[19]: 1) developing an ML model using predictive algorithms, 2) applying human validation and annotation with expert oversight, 3) conducting model training where input data is organized and refined, and 4) real-world deployment while considering ethical implications. Predictive algorithms such as Proximal Policy Optimization (PPO) [20] are initially used, with human experts monitoring their functionalities where they replace dangerous actions in critical situations and modify or supplement the reward function. Moreover, Active Learning (AL) can detect instances where AVs lack adequate confidence and refer them to human annotators [21]. The model is trained after human validation, annotation, and feedback from human experts. Meanwhile, Curriculum Learning (CL) with the knowledge of human experts provides sorted and organized input for AVs [22], facilitating and expediting the learning process. Finally, the model is deployed in the real world, considering ethical implications.
Rapid progress in user interface design [23], regulatory frameworks [24], AV security [25][26] and explainable AI [27] have significantly improved the practicality of HITL approaches for AVs. Improved user interfaces provide intuitive controls and real-time feedback, while regulatory requirements ensure human oversight for safety and ethical compliance. Developments in AV security safeguard against cyber threats, and research in explainable AI enhances the transparency of decision-making processes, making it easier for human operators to understand and interact with AVs. Together, these advancements enable effective human oversight and intervention, crucial for the safe and ethical operation of AVs.
@@ I-B Existing Surveys and Tutorials
In the literature, a few works consider HITL-ML. For example, Duan etal. [28] provide an up-to-date overview of the latest results on understanding and utilizing the human factor in different contexts of mobile networks. Expanding on HITL-ML strategies, Mosqueira-Rey etal. [29] present a comprehensive review of state-of-the-art techniques, highlighting different strategies for incorporating humans into the ML process. From a data perspective, Wu etal. [19] survey HITL-ML, addressing how human interaction can enhance data quality and model performance. Rajendran etal. [30] review the main concepts and methods for HITL learning, paving the way for developing frameworks that facilitate human-machine teaming through safe learning and anomaly prediction. Methodologies, challenges, and opportunities associated with real-world applications of HITL systems are reviewed by Kumar etal. [31], discussing key issues like data quality, bias, and user engagement. Wilchek etal. [32] review the literature on HITL, emphasizing the role of human interaction in enhancing computer vision models performance and accuracy.
Specifically, in the realm of CL within RL, Narvekar etal. [33] provide a framework and an overview of existing CL methods. They classify these methods based on their assumptions, capabilities, and goals, offering insights into how CL can systematically improve RL performance. Soori etal. [34] examine recent advances in AI, ML, and DL in advanced robotic systems, and discuss their applications in modifying robots. Ma etal. [35] review the usage of AI for AVs development, and elaborate on critical issues such as functional needs and challenges in implementing AI-driven solutions for AVs. From a computer vision system design perspective, Nunes etal. [36] present a review of HITL-Cyber-Physical Systems (CPSs), providing insights into how human factors can be integrated into CPSs to enhance their functionality and reliability. AD technologies have seen substantial advancements, driven by developments in ML and AI. Chib and Singh [37] review the End-to-End AD stack, categorizing automated driving tasks using neural networks in an End-to-End manner. Surveying essential components and cutting-edge technologies, Zhao etal. [38] explore the development and implementation of AD systems. This survey highlights the key technologies that enable AD and the challenges associated with their deployment. Both Wang etal. [39] and Moura etal. [40] address ethical dilemmas associated with different aspects of AVs.
@@ I-C Contribution
In this paper, we review HITL-ML to support the safe and ethical operation of AVs, facilitate the management of complex and unpredictable scenarios from busy local roads to intricate intersections, trajectories and missions, and improve their trustworthiness, enhancing public confidence on them. Our paper differs from others in the literature by emphasizing the critical role of HITL-ML in ensuring the safe and ethical operation of AVs. We focus on CL, HITL-RL and AL (see Table II for a comparative summary of recent research papers on these techniques) and then we look into ethical principles tailored to the unique challenges and opportunities in AVs.
The key contributions of this paper are listed as follows:
  • We provide a comprehensive overview of the application of CL in AVs to improve generalization and convergence speed by structuring the training process from simpler to more complex tasks. We highlight the synergy between CL and RL and how their combination can reduce training time and improve AV performance on tasks such as navigation, path planning and obstacle avoidance.
  • We provide an in-depth overview of the latest HITL-RL methods and their applications, highlighting how human involvement can enhance the RL process through techniques such as reward shaping, action injection and interactive learning. We also refer examples from recent studies showing various HITL-RL frameworks that utilize human feedback to refine AV learning processes and ensure real-time safety interventions.
@@ I-D Paper Structure and Organization
Section II introduces CL and its significant role in improving learning speed in AVs. Section III introduces HITL-RL and how humans can cooperate in the RL process in AVs. Section IV discusses AL and examines its role in facilitating data annotation in AVs. Section V addresses the ethical aspects of AVs and provides guidelines to achieve ethical AVs using HITL-ML. Section VI highlights open research directions and, finally, we conclude the paper in Section VII.
@@
[41] assist the UAV in obstacle avoidance. DRL Navigation / UAV
[42] improve the agent’s learning ability and performance. DRL Intelligence / AV
[43] tackle sparse reward and low learning efficiency. RL Exploration / AUV
[44] facilitate mapless obstacle avoidance. DRL Navigation / UAV swarm
[45] improve the learning performance of RL DRL AD
[46] mitigate in-flight UAV anomalies by using semi-supervised SVM AL Anomaly detection / UAV
[47] develop a new path planning framework for AL AL Path planning / Semantic mapping
[48] detect unknown network attacks in the UAV network. AL IDS / UAV
[21] facilitate manual labeling of large-scale surveillance data. AL VTR / ITS
[49] learn the decentralized flocking and obstacle avoidance policy CL Obstacle avoidance / UAV swarm
[50] reduce the required training time and improve agent performance CL AD
[51] enable mission planning for controlling a UAV ML Mission planning / UAV
[52] address 3D object detection in AV AL Object detection / UAV
[53] study collision-free flocking control problem for scalable and large-scale fixed-wing UAV CL Flocking problem / UAV swarm
[54] learn an end-to-end urban driving policy for CARLA driving simulator CL Driving simulator
@@ II Curriculum Learning
Figure 2: With CL for AVs tasks are divided into smaller, manageable parts, progressing from easy to more difficult challenges. Data associated with each part is sequentially fed to the model to enhance learning efficiency and improve overall performance.
@@
CL is an ML technique that structures the training process by starting with simpler tasks and gradually progressing to more complex ones. This approach enhances generalization and convergence speed, making it particularly effective for training AVs [55]. Recent studies have applied CL to AVs in different ways, including curriculum-based multi-agent DRL for decentralized flocking and obstacle avoidance, collision-free flocking control in large-scale UAV swarms, and adaptive curriculum strategies for efficient target search in disaster scenarios using UAV swarms[56]. These applications demonstrate the effectiveness of CL in guiding autonomous agents through progressively challenging tasks to achieve optimal performance.
@@
Fig. 2 illustrates the concept of CL for AVs. Tasks are divided into smaller, manageable parts, progressing from easy to more difficult challenges. This method allows the AVs to build foundational skills with simpler tasks before gradually tackling more complex scenarios, thereby improving learning efficiency and performance over time. The structured progression ensures that the vehicle can handle increasingly difficult situations with greater accuracy and reliability. In the realm of AVs, the combination of CL and RL can reduce the required training time, improve agent performance, and help learn better policies [50]. AVs require mastery of complex tasks such as navigation, path planning, and obstacle avoidance. CL can aid the learning process by ordering these tasks from simple to complex.
II-A Literature Review
II-A1 Navigation and Data Collection